OpenStudy (truredd):

Two objects are moving along separate linear paths where each path is described by position, d, and time, t. The variable d is measured in meters, and the variable t is measured in seconds. The equation describing the graph of the position of the first object with respect to time is d = 2.5t + 2.2. The graph of the position of the second object is a parallel line passing through (t = 0, d = 1). What is the equation of the second graph?

OpenStudy (truredd):

d = 2.5t + 1 d = t + 2.5 d = -0.4t + 1 d = 2.5t + 3.2

OpenStudy (welshfella):

Hint:- as the lines are parallel the graph will have the same slope as the line in the question remember the general formula y = mx + b , where m = the slope.

OpenStudy (welshfella):

so the slope of d = 2.5t + 2.2 = what?

OpenStudy (truredd):

a?

OpenStudy (truredd):

medal given

OpenStudy (truredd):

huh